OpenStudy (anonymous):

what is the area of a triangle with sides of 2 and 5?

OpenStudy (paki):

use formula... 1/2(b*h)

OpenStudy (anonymous):

sooo... (2x5)1/2?

OpenStudy (paki):

yeah,,, 1/2(2*5)= 10/2= 5.

OpenStudy (anonymous):

so the area is 5, even though one of the sides is 5?!?...

OpenStudy (paki):

yeah 5

OpenStudy (anonymous):

ok thanks

OpenStudy (paki):

pleasure